# Autoregressive Language Model Pre-training

### v1

@@ 4.2 Main Results
We evaluate trained models on language modeling evaluation sets. As shown in Figure 2 (right), our MoE models outperform the dense baseline in all domains, significantly reducing perplexity. For example, our 0.3B/32E model achieves a relative improvement of 13.9% on Books compared to the 0.3B dense model. We observe that the improvement is especially large in test domains that are markedly different from the domains of the training dataset (e.g. Python). We consider this as a strong indication of expert specialization in specific domains (We further study expert specialization in Section 5.4).
@@ 5 Analysis and Ablation Studies
In this section, we conduct ablation studies and analysis to understand the essence of each component of our approach.
@@ 5.2 Importance of Similarity-based Data Batching
5.3 Comparison with Existing MoE Models
We compare our approach with Expert Choice (EC) (Zhou et al., 2022), a state-of-the-art MoE method that ensures balanced load during training by having each expert select top-$k$ inputs according to the routing weights. We consider two variants of EC MoE models, both with a capacity factor of 1 to match the computation of our MoE models. First, we train a sparse EC MoE model using our segment routing strategy, where each expert selects top segments and processes all tokens within those segments. This variant allows us to directly compare our expert-merging strategy with the expert choice method while using the same segment-level routing approach.
@@ 5.3 Comparison with Existing MoE Models
Second, we consider the original EC setting with token-level routing to provide an end-to-end comparison with state-of-the-art MoE models using the same amount of training computation. Figure 5 shows the training loss curves. We observe that Lory (blue curve) significantly outperforms segment-level EC (orange curve) with the same routing setting, suggesting that a fully differentiable architecture is more effective than a sparse MoE when using the same routing strategy. Comparing Lory with the token-level EC model (green curve), we find that Lory achieves competitive results despite using segment-level routing and not requiring any advanced training techniques. These results highlight the significant potential of Lory. In Appendix G.1, we compare Lory and EC on held-out evaluation sets. We find Lory achieves much better perplexity compared to the token-level EC model, while performing similarly on other domains (arXiv, Books, Wiki, C4). Our analysis in Section 5.4 demonstrates that Lory learns experts specialized in specific domains (e.g., Python code), potentially improving performance in less frequent domains.
@@ 5.4 Expert Utilization and Specialization
5.5 More Analysis and Discussion
In Appendix G, we further show that (1) during inference of downstream tasks, routing the entire input prompt once or routing each segment does not make substantial differences on the tasks we evaluate; (2) warmup training is crucial to achieve high expert utilization, especially when training MoE models with a large number of experts. In addition, we discuss training parallelism strategies when further scaling up model sizes in Appendix H.1; and discuss the potential of converting Lory to sparse models for more efficient inference in Appendix H.2.
@@ 6 Related Work
Our approach enables end-to-end gradient back-propagation by employing fully differentiable MoE architectures. SMEAR (Muqeeth et al., 2023) proposes softly merging experts by taking a weighted average on the parameter space. However, SMEAR is only applied to text classification tasks with an encoder backbone. Although Lory shares a similar expert merging technique, it is the first approach that scales such architecture to autoregressive language model pre-training. Soft MoE (Puigcerver et al., 2024) is another fully-differentiable MoE architecture which enables end-to-end gradient back-propagation. However, it is only evaluated on vision tasks and does not apply to autoregressive language model pre-training either. We leave how to extend Soft MoE on decoder language models as the future work.
@@ 7 Conclusion
In this paper, we propose Lory, a fully-differentiable MoE model designed for autoregressive language model pre-training. Our extensive experiments demonstrate that Lory significantly outperforms its dense counterpart on language modeling and downstream tasks. We also observe that trained experts are highly specialized and capable of capturing domain-level information. Future research includes further scaling up Lory, combining token-level routing and segment-level routing, and developing efficient decoding methods for Lory.
@@ Acknowledgements
We appreciate useful comments and feedback from the members of the Princeton NLP group. We thank Weijia Shi for the help with the experiments and discussion related to the similarity-based data batching method.
@@
This paper presents a new approach for building large language models. We would like to note that, similar to existing language models, the language models trained with our approach may have the same potential societal consequences. For example, language models can produce factually inaccurate outputs, which carry the risk of spreading misinformation (e.g., (Min et al., 2023)). Additionally, malicious users can extract training data used to train language models, potentially causing privacy and licensing issues (Carlini et al., 2021). We acknowledge these potential negative consequences and caution those who use our approach to build powerful language models.
@@ Appendix B Computational Overhead of Routing and Merging
Here we investigate the computational overhead of our Moe models compared to the dense counterpart. We consider an MoE layer and an input tensor $x$ consisting of $L$ tokens and $d$ dimensions: $x:L\times d$. We assume that the model uses SwiGLU as the activation function in FFNs and it up-projects the input $x$ to $d^{\prime}$-dimensional activations in FFNs. In this case, processing the input on an FFN requires roughly $6\times L\times d\times d^{\prime}$ FLOPs (there are two up projections and one down projections in SwiGLU-based FFNs). The overhead of soft-routing MoE comes mainly from the merging operation. Suppose that there are $E$ experts and that the model makes a routing decision for every segment of $T$ tokens (equivalently, there are $L/T$ routing decisions). Each merging operation on $E$ experts takes $6\times E\times d\times d^{\prime}$ FLOPs (we compute three merged matrices). Therefore, the total overhead will be $\frac{L}{T}\times 6\times E\times d\times d^{\prime}$ FLOPs. This indicates that compared to a dense FFN layer, an MoE layer with $E$ experts requires $\frac{E}{T}$ more FLOPs, compared to the dense counterpart. In our experiments, we set $T=256$; this suggests that using $E=8$ experts introduces $3.1\%$ more computations and using $E=32$ experts introduces $12.5\%$ more computations at the FFN/MoE layers. It is worth noting that the computations from FFN layers are only a subset of the full model computations, so $3.1\%$ is an overhead upperbound when measuring on full models. In our experiments, our most straightforward implementation leads to a 15% or 28% slowdown of training efficiency when using 8 or 32 experts (Table 2). We discuss the training strategy for scaling up model sizes in Appendix H.1 and leave a better-optimized implementation for future work.
@@ Appendix G More Analysis and Ablation Studies
G.1 Held-out Evaluation of Lory and EC models
Model arXiv Books Wiki C4 Python
0.3B/8E (Lory) 7.4 16.0 9.2 13.3 12.5
0.3B/8E (EC, seg-level) 7.9 17.6 10.5 14.1 14.1
0.3B/8E (EC, tok-level) 7.5 17.0 9.2 12.8 13.6
Table 8 shows the perplexity of the models on held-out evaluation sets. We find that the token-level EC model performs better than our model on the C4 evaluation data, which is likely on the most similar distribution to the training set (Commoncrawl); on arXiv, Books, and Wikipedia, EC performs similarly or slightly worse than our model. Interestingly, on the Python evaluation set, our model achieves a much better perplexity compared to the token-level EC model (12.5 vs 13.6). We suspect that segment-level routing may be particularly effective at handling out-of-domain evaluation data (we assume that there is only a very small part of Python code in Commoncrawl). Our analysis in Section 5.4 shows that segment-level routing models are indeed able to learn experts that are specialized in specific domains (e.g., Python code), potentially helping models achieve high performance in less frequent domains.
@@ G.2 Expert Utilization
G.3 Inference Methods
@@
In addition, we discuss the possibility of converting Lory to sparse models for more efficient inference in Appendix H.2.
@@
G.4 Warmup Training
@@
Appendix H Discussion
H.1 Parallelism for Scaling
In our experiments, we use data parallelism with ZeRO optimization (Rajbhandari et al., 2020) to reduce memory usage. Data parallelism either saves a copy of the model parameters in each device or relies on fast communication between devices to synchronize parameters if using the ZeRO optimization. When we scale up the MoE models (e.g. model sizes $>100$ B), this introduces a potential problem as the amount of MoE parameters increases dramatically, resulting in a bottleneck due to significant parameter communication.
To address this problem, expert parallelism has been utilized for sparsely activated MoE models, where experts are distributed across different devices and host independently (Lepikhin et al., 2021; Fedus et al., 2022). Unlike traditional MoE models that involve a top-k discrete routing network to dispatch inputs to a subset of experts, our soft-routing MoE models require access to all expert parameters for the computations. To overcome this challenge, we employ an expert-wise model parallelism strategy. This involves partitioning the MoE layers across multiple devices by sharding all experts along with their hidden dimension. Consequently, same dimensions from different experts are allocated to the same device. As Figure 10 shows, we shard all the experts along with the hidden dimension, and distribute the same shard of different experts to one device. The expert-wise model parallelism enables us to communicate only hidden activations instead of model parameters. This is much more efficient as the hidden activations do not scale with the number of experts.
Furthermore, for non-MoE components (e.g., attention layers), we can still employ data parallelism to avoid communication overhead. Figure 11 illustrates a training setup in which we employ a mixture of data parallelism and model parallelism for training at scale.
Figure 10: Comparison between expert parallelism and expert-wise model parallelism. In soft-routing MoE models, the access to all expert parameters is needed for the computations. We shard all the experts along with hidden dimension, and distribute them across different devices.
Figure 11: Illustration of using a mixed strategy of data parallelism and expert-wise model parallelism. For non-MoE modules (e.g., attention layers), we use the data parallelism; while for MoE layers, we employ the model parallelism to avoid significant parameter communication.
H.2 Converting to Sparse Models for Inference

### v2

@@ 4.2 Main Results
We evaluated the models on language modeling tasks and found that our MoE models consistently outperformed the dense baseline, significantly reducing perplexity across all domains. For instance, the 0.3B/32E model improved perplexity by 13.9% on Books compared to the 0.3B dense model. Notably, improvements were most pronounced in test domains distinct from the training data (e.g., Python), indicating strong expert specialization, which is further explored in Section 5.4).
@@ 5 Analysis and Ablation Studies
In this section, we conduct ablation studies and analysis to understand the significance of each component of our approach. In Appendix G, we provide additional insights, demonstrating that (1) during the inference of downstream tasks, whether the entire input prompt is routed once or each segment is routed individually does not result in substantial differences in the tasks we evaluated; and (2) warmup training is crucial for achieving high expert utilization, particularly when training MoE models with a large number of experts.
@@ 5.2 Importance of Similarity-based Data Batching
5.3 Comparison to Existing MoE Models
We compare our approach with Expert Choice (EC) (Zhou et al., 2022), a state-of-the-art MoE method that ensures balanced load during training by having each expert select top-$k$ inputs according to the routing weights. During inference, we route each token into top-$k$ experts to avoid to leverage global information for routing.
@@ 5.3 Comparison to Existing MoE Models
We consider two variants of EC MoE models, both with a capacity factor of 1 to match the computation of our MoE models. First, we train a sparse EC MoE model using our segment routing strategy, where each expert selects top segments and processes all tokens within those segments. This variant allows us to directly compare our expert-merging strategy with the expert choice method while using the same segment-level routing approach. Second, we consider the original EC setting with token-level routing to provide an end-to-end comparison with state-of-the-art MoE models using the same amount of training computation. Figure 5 shows the training loss curves. We observe that Lory (blue curve) significantly outperforms segment-level EC (orange curve) with the same routing setting, suggesting that a fully differentiable architecture is more effective than a sparse MoE when using the same routing strategy. Comparing Lory with the token-level EC model (green curve), we find that Lory achieves competitive results despite using segment-level routing and not requiring any advanced training techniques. These results highlight the significant potential of Lory.
Table 2 shows model perplexity on held-out evaluation sets. The token-level EC model outperforms ours on C4, likely due to its similarity to the training set (Commoncrawl). However, on arXiv, Books, and Wikipedia, EC performs similarly or slightly worse. Notably, our model excels on the Python evaluation (12.5 vs. 13.6 perplexity), suggesting segment-level routing can be particularly effective for out-of-domain data (i.e., Python in CommonCrawl). Our analysis in Section 5.4 shows that segment-level routing models are indeed able to learn experts that are specialized in specific domains (e.g., Python code), potentially helping models achieve high performance in less frequent domains.
Model arXiv Books Wiki C4 Python
0.3B/8E (Lory) 7.4 16.0 9.2 13.3 12.5
0.3B/8E (EC, seg-level) 7.9 17.6 10.5 14.1 14.1
0.3B/8E (EC, tok-level) 7.5 17.0 9.2 12.8 13.6
@@ 5.4 Expert Utilization and Specialization
6 Discussion of Parallelism Strategies for Scaling Lory
In our experiments, we utilized data parallelism with ZeRO optimization (Rajbhandari et al., 2020) to reduce memory usage. Data parallelism involves either maintaining a copy of the model parameters on each device or using fast communication between devices to synchronize parameters with ZeRO optimization. As we scale up the Mixture of Experts (MoE) models to sizes exceeding 100 billion parameters, the substantial increase in MoE parameters can lead to a bottleneck due to the significant amount of parameter communication required.
To address this problem, expert parallelism is utilized for sparsely MoE models, where experts are distributed across different devices and operate independently (Lepikhin et al., 2021; Fedus et al., 2022). Unlike traditional MoE models, which use a top-k discrete routing network to dispatch inputs to a subset of experts, our soft-routing MoE models require access to all expert parameters for computations.
To overcome this challenge, we can employ an expert-wise model parallelism strategy, as shown in Figure 7 (left). This involves partitioning the MoE layers across multiple devices by sharding all experts along with their hidden dimension. Consequently, identical dimensions from different experts are allocated to the same device. We shard all the experts along with the hidden dimension and distribute the same shard of different experts to one device. This expert-wise model parallelism enables us to communicate only hidden activations instead of model parameters, which is more efficient since hidden activations do not scale with the number of experts. Furthermore, for non-MoE components, such as attention layers, we can employ data parallelism to minimize communication overhead. Figure 7 (right) illustrates a training setup in which we use a combination of data parallelism and model parallelism to facilitate training at scale. We plan to explore the experiments for scaling our approach to very large models in future work.
Figure 7: Left: Comparison between expert parallelism and expert-wise model parallelism. Right: Using a mixed strategy of data parallelism and expert-wise model parallelism. We use data parallelism for non-MoE components (e.g., attention) and model parallelism for MoE layers.
@@ 7 Related Work
Our approach enables end-to-end gradient back-propagation using fully differentiable MoE architectures. While SMEAR (Muqeeth et al., 2023) and Soft MoE (Puigcerver et al., 2024) also achieve this, SMEAR is limited to text classification tasks with an encoder backbone, and Soft MoE is only evaluated on vision tasks. Lory is the first to scale this architecture for autoregressive language model pre-training. Extending Soft MoE to decoder language models is left for future work.
@@ 8 Conclusion
In this paper, we introduce Lory, a fully differentiable MoE model specifically designed for autoregressive language model pre-training. Through extensive experiments, we demonstrate that Lory significantly outperforms its dense counterparts in both language modeling and downstream tasks. Our findings reveal that the trained experts in Lory are highly specialized and adept at capturing domain-level information. Future research directions include scaling up Lory further (Section 6), integrating token-level and segment-level routing, and developing efficient decoding methods tailored for Lory.
@@ Acknowledgements
We appreciate the valuable comments and feedback from members of the Princeton NLP group. We also thank Weijia Shi for the help with the experiments and discussions related to the similarity-based data batching method.
@@
This paper presents a novel approach for training sparse large language models. It is crucial to acknowledge that, similar to existing language models, those trained using our method may have comparable societal consequences. For instance, language models can generate factually inaccurate outputs and unethical content, posing risks of spreading misinformation and perpetuating harmful stereotypes or biases (Zhao et al., 2024). Furthermore, malicious users may extract the training data used to develop these models, leading to potential privacy and licensing issues (Carlini et al., 2021). We recognize these potential negative consequences and advise those employing our approach to remain vigilant about these risks when developing powerful language models.
@@ Appendix B Computational Overhead of Routing and Merging
Here we investigate the computational overhead of our Moe models compared to the dense counterpart. We consider an MoE layer and an input tensor $x$ consisting of $L$ tokens and $d$ dimensions: $x:L\times d$. We assume that the model uses SwiGLU as the activation function in FFNs and it up-projects the input $x$ to $d^{\prime}$-dimensional activations in FFNs. In this case, processing the input on an FFN requires roughly $6\times L\times d\times d^{\prime}$ FLOPs (there are two up projections and one down projections in SwiGLU-based FFNs). The overhead of soft-routing MoE comes mainly from the merging operation. Suppose that there are $E$ experts and that the model makes a routing decision for every segment of $T$ tokens (equivalently, there are $L/T$ routing decisions). Each merging operation on $E$ experts takes $6\times E\times d\times d^{\prime}$ FLOPs (we compute three merged matrices). Therefore, the total overhead will be $\frac{L}{T}\times 6\times E\times d\times d^{\prime}$ FLOPs. This indicates that compared to a dense FFN layer, an MoE layer with $E$ experts requires $\frac{E}{T}$ more FLOPs, compared to the dense counterpart. In our experiments, we set $T=256$; this suggests that using $E=8$ experts introduces $3.1\%$ more computations and using $E=32$ experts introduces $12.5\%$ more computations at the FFN/MoE layers. It is worth noting that the computations from FFN layers are only a subset of the full model computations, so $3.1\%$ is an overhead upperbound when measuring on full models. In our experiments, our most straightforward implementation leads to a 15% or 28% slowdown of training efficiency when using 8 or 32 experts (Table 3).
@@ G.1 Expert Utilization
G.2 Inference Methods
Comparison to segment-level routing during inference.
@@ G.2 Inference Methods
Converting to sparse models for efficient inference.
G.3 Warmup Training